OpenStudy (anonymous):

PLEASE HELP!!! If a baseball player hits a baseball from 4 feet off the ground with an initial velocity of 64 feet per second, how long will it take the baseball to hit the ground?

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

can you help me?

OpenStudy (anonymous):

what grade r u in

OpenStudy (anonymous):

10

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

i thought it was easy 4th grade math sorry i'm a 6th grader.

OpenStudy (anonymous):

um, okay

OpenStudy (anonymous):

sorry bye

OpenStudy (anonymous):

@Vocaloid

OpenStudy (ddcamp):

You can use the equation: \[y = y_{initial} + v_{initial}\times t - 16 \times t^2 \] to find the height of the ball after t seconds. Solve the equation for y=0 and you'll get the time where the height is 0 (i.e. its on the ground)

OpenStudy (anonymous):

How come the direction of acceleration is not specified? You would need to identify the horizontal velocity and vertical velocity separately to know when exactly the ball will hit the ground.

OpenStudy (anonymous):

so the answer is \[2 \pm \sqrt{17}/2 \] ?

OpenStudy (ddcamp):

@Robert136 The question doesn't ask where, just when.

OpenStudy (anonymous):

You don't know.

OpenStudy (ddcamp):

@RosieF The quadratic equation gives you two answers, because the graph of that function would look something like this: